### v2

# APVR: Hour-Level Long Video Understanding with Adaptive Pivot Visual Information Retrieval

Abstract
Current multimodal large language models (MLLMs) struggle with hour-level video understanding, facing significant challenges not only in modeling the substantial information volume of long videos but also in overcoming the memory wall and resource constraints during both training and inference. Although recent training-free approaches have alleviated resource demands by compressing visual features, their reliance on incomplete visual information limits the performance potential. To address these limitations, we propose Adaptive Pivot Visual information Retrieval (APVR), a training-free framework that hierarchically retrieves and retains sufficient and important visual information. It breakthroughs the memory wall limitation via two complementary components: Pivot Frame Retrieval employs query expansion and iterative spatio-semantic confidence scoring to identify relevant video frames, and Pivot Token Retrieval performs query-aware attention-driven token selection within up to 1024 pivot frames. This dual granularity approach enables the processing of hour-long videos while maintaining semantic fidelity. Experimental validations demonstrate significant performance improvements, achieving 64.9% on LongVideoBench and 68.4% on VideoMME, which are state-of-the-art results for both training-free and training-based approaches. Meanwhile, our method provides plug-and-play integration capability with existing MLLM architectures.
1 Introduction
The proliferation of long-form video content necessitates robust understanding capabilities that extend beyond the current limitations of multimodal large language models (MLLMs). While existing video-based models (Lin et al. 2023; Zhang et al. 2025) demonstrate proficiency on short video sequences, their scalability for hour-level video remains challenging. Specifically, they encounter not only the memory wall but also the inherent dilution of semantic information across extensive temporal spans.
The core challenge stems from the inability of current Video MLLMs to effectively model and extract critical information from long video sequences. In hour-level videos, semantically relevant contents are typically sparse across both temporal and spatial dimensions. However, existing approaches process entire sequences by uniform sampling, leading to computational inefficiency and degraded performance due to attention dispersion across irrelevant segments. As illustrated in Fig. 1(a), this strategy samples frames that are not pivot, leading to wrong answer generated by MLLMs.
Figure 1: Given a long video, uniform sampling (a) and insufficient frame retrieval (b) yield incorrect answer, while naively increasing pivot frames (c) raises OOM. Our APVR (d) explores joint frame-token co-retrieval, concentrating computational resources on the most relevant information.
Recently, some approaches (Ataallah et al. 2024; Tang et al. 2025a; Guo et al. 2025; Ye et al. 2025; Cheng et al. 2025; Luo et al. 2025) have attempted to retrieve query-relevant frames from long videos, simplifying long video understanding task as key-frame selection. Consequently, MLLMs process only few and sparse key frames instead of continuous video frames, neglecting the temporal-semantic relationships and potentially failing in scenarios requiring comprehensive temporal reasoning and long-range logical understanding, as shown in Fig. 1(b).
A potential solution to the above limitation is to retrieve sufficient pivot frames and improve coverage as shown in Fig. 1(c). However, such approaches encounter the memory wall when processing these frames with MLLMs. While recent training-based methods employ strategies such as multimodal sequence parallelism(Fu et al. 2025; Liu et al. 2024d) or feature compression (Shu et al. 2024; Liu et al. 2025a) during the training phase to accommodate extended video frames, they demand resource-intensive multi-stage training based on existing MLLMs. Thus, they lack adaptability to rapidly evolving MLLM architectures due to high retraining costs.
The inherent trade-off between temporal coverage and computational feasibility has hindered existing methods from achieving comprehensive video understanding. To address this core challenge, we propose Adaptive Pivot Visual information Retrieval (APVR) framework, a training-free approach that integrates spatio-semantic frame retrieval with token retrieval and compression, as shown in Fig. 1(d). This two-stage retrieval mechanism enables the system to concentrate computational resources on the pivot visual information that is most relevant to the input query, thus efficiently processing hour-long videos without encountering memory limitations.
Specifically, at the frame level, our Pivot Frame Retrieval (PFR) component expands the original query into four types of semantic information: objects, descriptions, relations, and semantics. This expansion enables more comprehensive frame scoring from spatial and semantic perspectives using complementary visual models: CLIP for semantic similarity and Grounding-DINO for object detection and spatial reasoning. The framework incorporates temporal diffusion mechanisms to maintain temporal coherence and employs adaptive resampling strategies to progressively refine frame selection across multiple iterations. At the token level, our Pivot Token Retrieval (PTR) component extends the concept of semantic importance to fine-grained visual representations. This component leverages query-aware multi-layer attention scoring to identify the most relevant visual tokens within selected frames, employing dynamic chunk-wise selection and head-wise soft voting mechanisms to maintain both computational efficiency and semantic accuracy.
The main contributions of this work are summarized as follows:
$\bullet$ We propose APVR, a novel training-free framework that addresses scalability challenges in long video understanding through dual-granularity information retrieval. Our approach combines frame-level pivot retrieval with token-level adaptive selection to breakthrough the memory wall.
$\bullet$ We develop a mechanism that leverages spatio-semantic confidence scoring and query-aware attention scoring to preserve temporal structure and semantic details, ensuring accurate understanding of complex video narratives while maintaining computational efficiency.
$\bullet$ We demonstrate that intelligent dual-granularity retrieval provides a sustainable alternative to parameter scaling for long video understanding. Our training-free design ensures plug-and-play integration with existing MLLM architectures while preserving foundational capabilities.
2 Related Work
Video MLLMs  Recent years have witnessed significant advances in multimodal large language models (MLLMs) for visual perception and understanding. As a critical modality in visual data analysis, video has also been integrated into MLLM frameworks. Notable efforts include Qwen2.5-VL (Bai et al. 2025a), InternVideo2.5 (Wang et al. 2025b), VideoLlama3 (Zhang et al. 2025), and VideoChat-Flash (Li et al. 2024b), which explore video understanding through a unified architecture consisting of three key components: (a) a video encoder for spatio-temporal feature extraction, (b) a projector for aligning visual features with linguistic embeddings, and (c) a video-aware LLM for multimodal reasoning. Due to the inherent complexity of video content (e.g., multi-object interactions, long-range temporal dependencies), existing MLLMs often fail to extract semantically salient information for video-centric downstream tasks requiring precise spatiotemporal reasoning, such as video question answering and spatiotemporal video grounding. Some works(Bai et al. 2025a; Wei et al. 2025) exploit to aggregate timestamps in the RoPE(Su et al. 2024). However, these methods exhibit significant performance degradation when extrapolating to videos exceeding their pre-training duration. Moreover, current architectures also exhibit limitations of inefficient cross-modal alignment between visual information and linguistic queries, particularly in long-form videos exceeding 5 minutes (Li et al. 2024a).
Long Video Understanding  Recently advances in long video understanding can be broadly categorized into training-based and training-free paradigms. In the training-based paradigm, MLLMs are directly trained using large-scale long video datasets. The whole training process is usually divided into several stages to stably improve different components of MLLM, as well as to improve the ability to model different types of input visual data (Li et al. 2024b; Zhang et al. 2025). In order to process extra-long videos, current methods explore two different strategies: sequence parallelism (Chen et al. 2024; Zhang et al. 2024b; Shen et al. 2025) or visual feature compression (Song et al. 2024; Shen et al. 2024; Liu et al. 2025a). For instance, LongVILA (Chen et al. 2024) adopts a five-stage training pipeline with a novel multi-modal sequence parallelism to efficiently understand long videos. Video-XL-Pro (Liu et al. 2025a) builds a learnable module to generate compact and comprehensive video tokens for hour-long video understanding. In the training-free paradigm, long video understanding task is organized as an agent-like process (Liu et al. 2025b; Wang et al. 2024c), where MLLMs serve as reasoning engine within the system. These training-free methods mainly focus on retrieving the most important information in video frames (Tang et al. 2025b; Park et al. 2024; Ye et al. 2025; Guo et al. 2025; Cheng et al. 2025) or visual tokens (Wang et al. 2025a; Luo et al. 2025). For example, AKS (Tang et al. 2025b) adopts a plug-and-play module for keyframe selection and information pre-filtering for video-based MLLMs. QuoTA (Luo et al. 2025) adopts an ante-hoc training-free modular for query-aware visual token assignment and important information retrieval. Despite these advancements, few methods simultaneously explore frame and token optimization, which is addressed in this work through adaptive pivot information retrieval.
3 Method
Figure 2: The overall framework of our proposed APVR. We integrate two plug-and-play components, Pivot Frame Retrieval and Pivot Token Retrieval, into MLLMs to improve the performance of long video understanding. APVR is training-free and provides a alternative to parameter scaling. With frame-level and token-level adaptive selection, it can accurately understand complex videos with computational efficiency.
3.1 Overview
The framework of the proposed APVR is illustrated in Fig. 2, which is a training-free system. Given an input video =$\mathcal{F}=\{f_{t}\}_{t=1}^{N}$ with $N$ frames sampled under specific $fps$ and query $Q$, the system aims to retrieve the most important visual information for accurate response. To achieve this, we build two main components in APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTF). In PFR, we search pivot frames $\mathcal{F}_{pivot}$ with high confidence scores using expanded queries and basic visual models. In PTR, we explore an attention-driven token selection in MLLMs to retrieve pivot tokens $\mathcal{T}_{pivot}$ that are most relevant with expanded queries.
3.2 Pivot Frame Retrieval
The progress of PFR is developed in an iterative and efficient manner, as elaborated in Alg. 1.
Semantic Information Expansion
First, We expand the given query $Q$ to four types of semantic information as follows:
Objects: The key objects $Obj_{k}$ or cue objects $Obj_{c}$ that are detectable by visual models such as Grounding-DINO (e.g., “bold person”,“phone”,“blue background”’). All objects are aggregated as $Obj=\{Obj_{k},Obj_{c}\}$.
Descriptions: The entities or hypernym concepts of objects based on knowledge graph augmentation, denoted as $Des$. (e.g. “pearl headband: decorative accessory worn on head; jewelry”)
Relations: The triplet relationship $Rel\subseteq Obj\times\mathcal{R}\times Obj$ to enhance the confidence of frames with logical links between objects. In each triplet $(o_{i},r,o_{j})$, the type of relation $r\in\mathcal{R}$ is exactly one of the following: (1) spatial which means that $o_{j}$ and $o_{j}$ appear in the same frame, such as “a bold man holding a red phone”. (2) time which means that $o_{j}$ and $o_{j}$ appear in different frames in orders such as “A woman walks through, after a cat follows.” (3) attribute which means that $o_{i}$ appears with an attribute describable by $o_{j}$ such as “A woman in a white skirt”. (4) causal which means that $o_{j}$ and $o_{j}$ appear in orders with a cause-effect manner such as “A woman opens the door, the cat is scared and runs out.”
Semantics: The semantics information of query and options based on knowledge graph, denoted as $Sem$ (e.g., ‘leash often appears with dog”).
The prompt template for LLM-based query expansion is provided in Appendix.
Algorithm 1 Pivot Frame Retrieval (PFR)
1:video frames $\mathcal{F}$, query $Q$, iterations $P$, stride $\nabla$
2:Pivot frames $\mathcal{F}_{pivot}=\{f_{t}\}_{t=1}^{K}$
3:Initialize:
4:$\mathcal{S}\leftarrow\mathbf{0}^{N}$triangleright Confidence scores
5:$\mathcal{V}\leftarrow\emptyset$triangleright Visited frames set
6:$\{Obj,Des,Rel,Sem\}\leftarrow\operatorname{LLM}(Q)$triangleright Query Expansion
7:for $p=1$ to $P$ do
8:     $\nabla_{p}\leftarrow\max\left(1,\frac{\nabla}{p}\right)$
9:     $\mathcal{F}_{samp}^{p}\leftarrow\operatorname{AdaptiveResample}(\mathcal{F},\mathcal{S},\mathcal{V},\nabla_{p})$
10:     for $f_{t}\in\mathcal{F}_{samp}^{p}$ do
11:         $s^{CLIP}_{t}\leftarrow\Pi_{CLIP}(f_{t},Q,Des,Sem)$
12:         $s^{GD}_{t}\leftarrow\Pi_{GD}(f_{t},Obj,Rel)$
13:         $\mathcal{S}_{t}\leftarrow(1-\lambda)s^{CLIP}_{t}+\lambda s^{GD}_{t}$
14:         $\mathcal{S}_{[t-w,t+w]}\leftarrow\max\left(\mathcal{S}_{[t-w,t+w]},\frac{\mathcal{S}_{t}}{1+|i-t|}\right)$
15:     end for
16:     $\mathcal{V}\leftarrow\mathcal{V}\cup\mathcal{F}_{samp}^{p}$
17:end for
18:return $\mathcal{F}_{pivot}\leftarrow\operatorname{topk}(\mathcal{S}\odot\mathcal{F},K)$
Spatio-Semantic Confidence Scoring
Given a video $\mathcal{F}$, and query $Q$ with expanded semantic information $\{Obj,Des,Rel,Sem\}$, instead of exhaustively processing all $N$ frames, we uniformly sample few frames with initial stride $\nabla$ in the first iteration and adaptively resample frames for the subsequent iteration based on the online updated frame confidence. To ensure progressively finer sampling, we gradually decrease the stride by $\nabla^{p}=\mathrm{max}(1,\frac{\nabla}{p}$). The frames sampled in iteration $p$ are denoted as $\mathcal{F}_{samp}^{p}$.
Our algorithm exploits two basic visual models, i.e., Grounding-DINO and CLIP, to comprehensively score $\mathcal{F}_{samp}^{p}$ as follows:
CLIP-based Similarity Scoring: We leverage the CLIP model to compute cross-modal similarity between text and visual features as the semantic confidence score based on expanded query and the visual feature by $s_{t}^{CLIP}=\Pi_{CLIP}(f_{t},Q,Des,Sem)$. The details of the function $\Pi_{CLIP}$ are provided in Appendix.
Grounding-DINO Object Detection: We leverage the Grounding-DINO model to compute the spatial confidence score based on the query-aware objects long with their relations by $s_{t}^{GD}=\Pi_{GD}(f_{t},Obj,Rel)$. The details of the function $\Pi_{GD}$ are provided in Appendix.
The final confidence score of this iteration is computed as a weighted sum of the CLIP and Grounding-DINO scores:
$\mathcal{S}_{t}=(1-\lambda)\cdot s_{t}^{CLIP}+\lambda\cdot s_{t}^{GD}.$ (1)
The frames scored in the current iteration are put into the visited frame set $\mathcal{V}$.
Temporal Diffusion: To account for the temporal continuity of video content and improve frame resampling, we introduce a temporal score diffusion mechanism that propagates confidence scores across neighboring frames as:
$\mathcal{S}_{i}=\mathrm{max}(\mathcal{S}_{i},\frac{\mathcal{S}_{t}}{1+|i-t|}),i\in[t-w,t+w]$ (2)
This diffusion process helps capture temporally coherent segments relevant to the query.
Adaptive Resampling
Based on the updated $\mathcal{S}$, we further design an adaptive and hybrid candidate sampling strategy for subsequent iterations that combines two types of set:
High-confidence Set: We select frames with higher scores from all unvisited frame set $\mathcal{U}=\{f_{t}\mid f_{t}\notin\mathcal{V}\}$ based on the distribution as:
$\mathcal{H}_{s}=\mathrm{TopK}(\mathcal{S}\odot\mathcal{U},\frac{N}{2\nabla_{p}})$ (3)
Uncertainty Set: We calculate Shannon entropy based on the score distribution as:
$\mathcal{E}_{\gamma}(s_{i})=-\sum_{j=i-\gamma}^{i+\gamma}e_{j}\cdot\text{log}(e_{j}),e_{j}=\frac{s_{j}}{\sum_{k=j-\gamma}^{j+\gamma}s_{k}}$ (4)
, then the frames with high entropy are identified by:
$\mathcal{H}_{e}=\{f_{t}\mid\mathcal{E}_{\gamma}(s_{t})>\mu+0.5\sigma\}$ (5)
, where $\mu$ and $\sigma$ are the mean and standard deviation of the entropy in $\mathcal{C}$.
The union candidate set for this iteration sampling is $\mathcal{C}=\mathcal{H}_{s}\cup\mathcal{H}_{e}$. Assume that there are $N_{\mathcal{C}}$ frames in $\mathcal{C}$. The frames $\mathcal{F}_{samp}^{p}$ are sampled in the way as:
$\mathcal{F}_{samp}^{p}=\mathcal{C}_{\text{multi}}\cup\mathcal{C}_{\text{rand}},$ (6)
$\mathcal{C}_{\text{multi}}\sim\mathrm{Multinomial}(S\odot\mathcal{C},(1-\alpha)N_{\mathcal{C}}),$ (7)
$\mathcal{C}_{\text{rand}}\sim\mathrm{Unif}(\mathrm{Perm}(S\odot\mathcal{C},\alpha N_{\mathcal{C}}))$ (8)
, where $\alpha$ balances the multinomial score-based importance sampling $\mathcal{C}_{\text{multi}}$ and the diverse random sampling $\mathcal{C}_{\text{rand}}$ to capture the full semantic content of the video and also to avoid falling into local optimal.
After $P$ iterations, the $K$ pivot frames are retrieved by $\mathcal{F}_{pivot}=\mathrm{TopK}(\mathcal{S}\odot\mathcal{F},K)$ and fed into the next module.
Model Params Type LongVideoBench VideoMME
Val Long Medium Short Overall
Proprietary Models
GPT4-V - training-based 59.1 53.5 - - 59.9
GPT4-o - training-based 66.7 65.3 - - 71.9
Gemini-1.5-Pro - training-based 64.0 67.4 - - 75.0
Open-Sourse MLLMs
VideoLLaVA(Lin et al. 2023) 7B training-based 58.2 36.2 38.0 45.3 39.9
VITA-1.5(Fu et al. 2025) 7B training-based - 47.1 54.2 67.0 56.1
mPLUG-Owl3(Ye et al. 2024) 7B training-based 52.1 50.1 57.7 70.0 59.3
VideoLLaMA3(Zhang et al. 2025) 7B training-based 59.8 54.9 63.7 80.1 66.2
Oryx-1.5(Liu et al. 2024c) 7B training-based 56.3 - - - 58.3
QuoTA(Luo et al. 2025) 7B training-free 59.0 55.7 64.9 77.1 65.9
Video-XL(Shu et al. 2024) 7B training-based 49.5 - - - 55.5
ViLaMP(Cheng et al. 2025) 7B training-based 57.8 - - - 67.5
AdaReTaKe(Wang et al. 2025a) 7B traning-free 62.6 58.3 - - 67.7
AKS(Tang et al. 2025a) 7B training-free 62.7 - - - 65.3
Kangaroo(Liu et al. 2024a) 8B traning-based 54.8 - - - 46.6
NVILA(Liu et al. 2024d) 8B training-based 57.7 54.8 62.2 75.7 64.2
ByteVideoLLM(Wang et al. 2024a) 14B training-based - 56.4 62.9 74.4 64.6
Qwen2-VL(Wang et al. 2024b) 7B training-based 55.6 53.8 - - 63.3
Qwen2-VL w/ APVR(ours) 7B training-free 60.9 55.1 66.2 73.9 65.2
Qwen2.5-VL(Bai et al. 2025b) 7B training-based 59.5 55.6 - - 65.4
Qwen2.5-VL w/ APVR(ours) 7B training-free 64.9 59.1 67.8 78.3 68.4
Table 1: Video understanding accuracy(%) on LongVideoBench and VideoMME. The proposed APVR is applied upon two baseline MLLMs. The Methods are divided into two categories: training-based and training-free. APVR based on Qwen2.5-VL achieved state-of-the-art results on both the two benchmarks.
Figure 3: Qualitative Comparison of APVR with the baseline MLLM. The expanded query is significant complementary for pivot information retrieval. The number and score of the pivot frame is drawn in the left-top with green and red text, respectively. Yellow stars indicate the key frame for correct response. Green answer represents correct while red one represents wrong.
3.3 Pivot Token Retrieval
Pivot Token Retrieval extends the concept of semantic importance from frame-level to token-level granularity. In this module, we adaptively retrieve the pivot visual token within MLLM to maintain computational efficiency while preserving critical semantic information.
First, the retrieved pivot frames $\mathcal{F}_{pivot}$ are processed by a Vision Encoder $\mathcal{G}(\cdot)$ and a Projector $\mathcal{P}(\cdot)$ to generate dense visual token representations: $\mathcal{T}_{vis}=\mathcal{G}(\mathcal{P}(\mathcal{F}_{pivot}))=\{\mathcal{T}_{t}\}_{t=1}^{K}$. Our objective is to retrieve the pivot token $\mathcal{T}_{pivot}$. To this end, we introduce two components: query-aware multi-layer attention scoring and adaptive token selection.
Query-aware Multi-Layer Attention Scoring
We leverage the expanded semantic information from the LLM-augmented query to guide token selection across multiple transformer layers. First, the attention scores for layer $l$ are computed as:
$A_{cross}^{l}=\mathrm{softmax}(\frac{q_{text}^{l}\cdot(k_{vis}^{l})^{T}}{\sqrt{d}})$ (9)
, where $A_{cross}^{l}\in R^{h\times d\times d_{q}\times d_{v}}$, $q_{text}^{l}$ denotes the query states derived from $\{Q,Des,Sem\}$ and $k_{vis}^{l}$ denotes the key state derived from the visual tokens $\mathcal{T}_{vis}$. Then the token-level attention scores of each layer are aggregated at the query dimension by:
$a^{l}=\sum_{d_{q}}A_{cross}^{l}.$ (10)
This formulation captures how much attention each key token receives across all query positions, providing a comprehensive measure of its importance for the current layer’s computations. The superscript $l$ will be omitted for concision.
Adaptive Token Selection
The simplest token selection strategy is to directly retain the Top-K tokens while evicting others. However, this approach ignores the continuity of visual tokens and may reduce accuracy. Another factor that matters is that attention scores across heads may exist disparities, which further result in independence between heads. Thus, considering the continuity of visual tokens as well as the disparities across heads, we introduce an adaptive token selection strategy consisting of two components: dynamic chunk-wise selection and head-wise soft voting.
Dynamic Chunk-wise Selection: The long video understanding reaches the memory wall when prefilling the visual token. Thus, we perform token selection for the KV cache to reduce redundancy. Given a key or value cache to select, we first divide it into $W$ chunks at the token dimension. Then, the base selection ratio of the chunk $w$ is computed as:
$\eta_{w}=\frac{\sum_{w}a}{\mathrm{max}(\{\sum_{i}a\mid i=1,...,W\})}$ (11)
The dynamic ratio of each chunk is calculated by:
$\rho_{w}=\mathrm{min}(1.0,\sqrt{\frac{|\{j:a_{j}>0.01\cdot\mathrm{max}(a)\}|}{L_{w}}})$ (12)
, where $L_{w}$ is the chunk length. The final selection ratio is determined by:
$\gamma_{w}=\rho_{w}\cdot\eta_{w}$ (13)
Head-wise Soft Voting Based on the ratio $\gamma_{w}$, we can select the pivot token in each chunk by:
$\mathcal{Z}_{w}=\mathrm{TopK}(a_{w}\odot\mathcal{T}_{w},\gamma_{w}L_{w})$ (14)
This results may appear to be significant disparities in different attention heads. To figure out this, we introduce head-wise voting for final token selection:
$\mathcal{Z}_{w}=\mathrm{TopK}((\sum_{j=1}^{h}\mathrm{softmax}(a_{w,j}))\odot\mathcal{T}_{w},\gamma_{w}L_{w})$ (15)
, where the softmax function calculates the per-head scores and gives a soft voting via normalization and sum. Finally, we update the KV cache as the selected pivot token in each layer:
$\mathcal{T}_{pivot}=\mathrm{Concat}(\{\mathcal{Z}_{w}\}_{w=1}^{W})$ (16)
Finally, the selected pivot tokens will be fed into the MLLM to generate the final response.
4 Experiments
4.1 Experimental setup
Datasets
We evaluate the performance of APVR in two popular benchmarks that contains hour-level videos: LongVideoBench(Wu et al. 2024). A benchmark for long-context video understanding, consisting of 3,763 videos with 6,678 annotated multiple-choice questions across 17 categories of referred reasoning questions. VideoMME(Fu et al. 2024). A multi-modal evaluation benchmark focused on fine-grained video understanding tasks requiring precise visual-textual alignment and detailed semantic comprehension. It comprises 900 videos totaling 256 hours, with 2,700 manually labeled complex multiple-choice question-answer pairs across 30 subfields. To highlight the importance of the pivot information retrieval for visual understanding, we do not use video subtitles for the two benchmarks.
Implementation Details
We integrate APVR with two baseline MLLMs, namely, Qwen2-VL-7B (Wang et al. 2024b) and Qwen2.5-VL-7B (Bai et al. 2025b). The MLLMs process a prompt involving the question, retrieved pivot video frames, and the multi-choice answer options to generate responses. We first densely extracted the raw frames from video with $fps=2$. Then, the sampled frames are iteratively scored using two basic visual models: CLIP (Radford et al. 2021) with ViT-B-16 (Dosovitskiy et al. 2020) backbone, and Grounding-DINO (Liu et al. 2024b) with Swin-T (Liu et al. 2021) backbone, based on the expanded semantic information generated by LLM. All evaluations are conducted on 8 NVIDIA A800 GPUs with 80GB memory, using LMMs-Eval (Zhang et al. 2024a) framework. Under the above setup, our APVR enables processing of up to 1024 frames, compared to the baseline 7B model’s capacity of 256 frames under optimal MLLM configurations.
4.2 Comparison to the State-of-the-Art
Quantitative results
We first compare the accuracy of video question answering between APVR with recent training-based or training-free methods. All results are summarized in Table 1. Upon the Qwen2-VL baseline, APVR improves the overall accuracy by 9.5% and 3.0% in LongVideoBench and VideoMME, respectively. Upon a stronger baseline Qwen2.5-VL, our APVR also improves the overall accuracy by 8.3% in LongVideoBench and 4.6% in VideoMME, respectively. Notably, APVR based on Qwen2.5-Vl reaches 64.9% accuracy in LongVideoBench, which is higher than the powerful proprietary models such as GPT4-V and Gemini-1.5-Pro. The overall performance of APVR also surpasses other training-based and training-free methods with larger parameter scales such as 8B and 14B. Another notable result is that our APVR reaches highest accuracy on long and medium video samples of VideoMME, confirming its unique capability to preserve critical spatio-temporal information in extended sequences.
Figure 4: The result of ablation study on LongVideoBench for the two designed components: PFR and PTR.
Qualitative results
Fig. 3 presents representative qualitative results on hour-level video understanding, comparing our APVR framework with the baseline Qwen2.5-VL across four challenging samples. The results clearly demonstrate the unique advantages of APVR in both the semantic expansion and the subsequent retrieval process.
Firstly, the semantic information expansion module in APVR significantly enriches the original query, providing the downstream retrieval process with comprehensive object, attribute, and relational cues that are otherwise missing when using only the raw query. This richer semantic context enables the retrieval module to more accurately identify and prioritize frames that are truly relevant to the query intent.
Secondly, we observe that the pivot frames selected by APVR are highly correlated with both the query and the expanded semantic information, effectively filtering out irrelevant or redundant content. In long videos, there often exist temporally adjacent frames that are visually similar and all somewhat related to the query. If only a very limited number of frames are selected, it is easy to miss critical moments; conversely, retaining an excessive number of frames (e.g., 1024 frames) would cause out-of-memory (OOM) errors. APVR overcomes this dilemma by PFR and PTR and distills the most informative visual information within sufficient pivot frames. As a result, APVR empowers MLLM to efficiently comprehend long-form video content, providing accurate and detailed answers even in scenarios where baselines fail or generate hallucinations.
4.3 Ablation Studies
Ablation studies on PFR and PTR
To systematically validate the effectiveness of the proposed dual-granularity retrieval strategy, we conduct ablation studies on the two core components of APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTR), using the LongVideoBench benchmark. The results, shown in Fig. 4, consistently demonstrate that removing either PFR or PTR leads to a notable decline in accuracy across both baseline MLLMs. Without PFR, the system loses its ability to precisely localize query-relevant frames, resulting in performance degradation due to the inclusion of irrelevant or redundant content. Without PTR, the model is less able to compress visual tokens and focus on the most informative visual details, leading to processing maximum to only 256 frames and reduced answer quality. These findings strongly support the necessity and rationality of our dual-granularity pivot information retrieval design.
Ablation studies on other hyper-parameters
We further conduct a comprehensive analysis of key hyper-parameters that influence the performance of APVR. First, we examine the impact of the number of retrieved pivot frames $K$ and the weighting parameter $\lambda$ that balances the contributions of the two visual models in frame scoring. As summarized in Table 2, increasing the number of retained frames generally improves performance, especially on long videos, by reducing the risk of missing relevant segments. For the model weighting parameter $\lambda$, we observe that optimal results are achieved by striking a balance between CLIP-based semantic similarity and Grounding-DINO-based object detection, confirming the complementarity of these two perspectives.
Second, we analyze the effect of core retrieval hyper-parameters, specifically the number of retrieval iterations $p$ and the initial sampling stride $\nabla$. As shown in Table 3, the best performance is observed when $P=3$ and $\nabla=4$, indicating that a moderate number of adaptive iterations and a balanced initial sampling density allow the system to progressively refine frame selection without causing overfitting or excessive computational cost. Deviating from these settings leads to decreased accuracy, highlighting the importance of carefully tuning the retrieval process to maintain a balance between sampling and distribution update.
$K$ Long Average $\lambda$ Long Average
1024 58.3 64.9 0.1 57.4 64.2
256 56.4 63.1 0.2 57.8 64.5
128 53.0 61.1 0.5 58.3 64.9
64 49.8 58.6 0.8 58.0 64.8
32 45.4 54.3 0.9 58.9 64.5
Table 2: Ablations of retained frame number and weight for scoring models.
$P$nabla Long Medium Short Average
3 2 56.4 62.1 77.3 62.9
3 4 58.3 65.8 77.3 64.9
3 8 55.1 64.6 77.3 63.1
2 4 56.0 66.0 77.3 63.9
4 4 58.0 65.8 77.3 64.7
5 4 58.5 64.5 77.3 64.5
Table 3: Ablations of iteration number and the initial sample stride.
5 Conclusion
We propose APVR, a training-free framework that overcomes the memory wall in hour-level video understanding via adaptive, hierarchical visual information retrieval. By integrating Pivot Frame Retrieval and Pivot Token Retrieval, APVR enables efficient and accurate processing of long videos without sacrificing semantic detail. Extensive experiments on standard benchmarks demonstrate that APVR delivers state-of-the-art results among training-free and training-based methods. Its plug-and-play design allows seamless integration with existing MLLM systems, eliminating the need for costly retraining. Our results establish retrieval-based optimization as a practical and scalable solution for long-form video comprehension.